OpenStudy (anonymous):

make x the subject of: y=-3 - 6/(x+2)

OpenStudy (campbell_st):

add 3 to both sides \[y + 3 = -\frac{6}{x+2}\] multiply both sides by (x+2) \[(x + 2)(y+3)= -6\] divide both sides by (y + 3) \[x + 2=-\frac{6}{y+3}\] subtract 2 from both sides \[x = -2 -\frac{6}{y+3}\]

OpenStudy (anonymous):

Thank you